OpenStudy (anonymous):

Hydrogen gas will give four strong spectral lines: Violet (400 nm), blue (425nm), cyan (476nm) and red (649nm). Calculate the energy associated with a photon of light at each of these wavelengths. Which has the highest energy? Which has the highest frequency? How do these compare relative to the respective wavelengths?

OpenStudy (jfraser):

use the 2 equations for wavelength, frequency, and energy of light\[c = \lambda * \nu = 3.0*10^8\frac{m}{s}\]and\[E = h*\nu\]where h is Planck's constant, \(6.626*10^{-34}J*s\)